# Coulomb excitation of hydrogen atoms by vortex ion beams

### v1

@@ IV Summary
In this work we have theoretically studied the Coulomb excitation of one–electron atoms by Laguerre–Gaussian ionic wave packets. Special attention was paid to the question of how the population of the magnetic sublevels of the excited atom is affected by the OAM projection of the projectile ions. In order to explore this OAM–dependence, we have applied the non–relativistic first–Born approximation and the density matrix theory. The general expressions were derived for the total excitation probability and the alignment parameter of the excited atom. Based on the theory obtained, detailed calculations have been carried out for the $\mathrm{H}(1s)+\mathrm{H}^{+}_{LG}\rightarrow\mathrm{H}(2p)+\mathrm{H}^{+}_{LG}$ excitation of hydrogen atoms by vortex proton beam. In these calculations, we have addressed not only a rather academic case of the excitation of a single well–localized atom, but also the much more promising—from the experimental viewpoint—scenario of collision of the ion beam with the macroscopic hydrogen target. For the latter case, particular emphasis was placed on the angular distribution of the subsequent $2p\to 1s$ radiative emission from the whole target, which is defined by the magnetic sublevel population of excited atoms and which can be easily measured in modern experiments. We have clearly shown that the anisotropy of the $2p\to 1s$ radiation and, hence, the alignment of the $2p$ state are very sensitive to the OAM projection of the Laguerre–Gaussian beam. This sensitivity opens us a very promising avenue for the diagnostics of vortex ion beams in accelerator and storage ring experiments
Acknowledgements
@@
A.S. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967.
@@
$\displaystyle\Psi^{LG}({\bf R},t)=\sqrt{\frac{n!}{(n+|\ell_{OAM}|)!}}\,\frac{1}{\pi^{3/4}}\left(\frac{R_{\perp}}{\sigma_{\perp}(t)}\right)^{|\ell_{OAM}|}\frac{1}{\sigma_{\perp}(t)\sqrt{\sigma_{z}(t)}}\,L_{n}^{|\ell_{OAM}|}\left(\frac{R_{\perp}^{2}}{\sigma_{\perp}^{2}(t)}\right)$ (15)
×$\displaystyle\times\exp\Big{\{}-\frac{i}{\hbar}t\frac{\langle p\rangle^{2}}{2m_{p}}+\frac{i}{\hbar}\langle p\rangle R_{z}+\frac{i}{\hbar}\ell_{OAM}\phi_{r}-i(2n+|\ell_{OAM}|+1)\arctan(t/t_{d,\perp})-\frac{i}{2}\arctan(t/t_{d,z})-$ (16)
$\displaystyle-\frac{R_{\perp}^{2}}{2\sigma_{\perp}^{2}(t)}\,(1-it/t_{d,\perp})-\frac{(R_{z}-v_{p}t)^{2}}{2\sigma_{z}^{2}(t)}\,(1-it/t_{d,z})\Big{\}}\,,$ (17)
@@
The set of solutions (18) for $n=0,1,2...$ and $l_{OAM}=0,1,2...$ is complete and orthogonal. It describes, moreover, free rectilinear motion on average,
@@
$t_{d,{\perp}}=\frac{m_{p}\hbar}{(\sigma^{p}_{\perp})^{2}}=\frac{\sigma^{2}_{\perp}(t_{0})}{c\lambda_{c}}$ (22)
is the time after which $\sigma_{\perp}(t_{d})=\sqrt{2}\sigma_{\perp}(t_{0})$, $\lambda_{c}=\hbar/m_{p}c$ is the Compton wavelength of the proton, and $\sigma^{p}_{\perp}$ is the width of the wave packet in the momentum space. Similar expressions can be obtained for the longitudinal direction:
@@
$\displaystyle\sigma_{z}(t_{0})=\frac{\hbar}{\sigma^{p}_{z}},\ t_{d,z}=\frac{m_{p}\hbar}{(\sigma^{p}_{z})^{2}}=\frac{\sigma^{2}_{z}(t_{0})}{c\lambda_{c}}.$ (24)
One can note that for both cases $t_{d,{\perp}}$ and $t_{d,z}$ are inversely proportional to $c\lambda_{c}=h/m_{p}\approx 4\times 10^{-7}$ m2/s, which implies characteristic dispersion time of about $t_{d}\sim 10^{-6}$ s for $\mu$m–size proton wave packets.

### v2

@@ 4 Summary
In this work we have theoretically studied the Coulomb excitation of one–electron atoms by Laguerre–Gaussian ionic wave packets. Special attention was paid to the question of how the population of the magnetic sublevels of the excited atom is affected by the OAM projection of the projectile ions. In order to explore this OAM–dependence, we have applied the non–relativistic first–Born approximation and the density matrix theory. The general expressions were derived for the total excitation probability and the alignment parameter of the excited atom. Based on the theory obtained, detailed calculations have been carried out for the $\mathrm{H}(1s)+\mathrm{H}^{+}_{LG}\rightarrow\mathrm{H}(2p)+\mathrm{H}^{+}_{LG}$ excitation of hydrogen atoms by vortex proton beam. In these calculations, we have addressed not only a rather academic case of the excitation of a single well–localized atom, but also the much more promising—from the experimental viewpoint—scenario of collision of the ion beam with the macroscopic hydrogen target. For the latter case, particular emphasis was placed on the angular distribution of the subsequent $2p\to 1s$ radiative emission from the whole target, which is defined by the magnetic sublevel population of excited atoms and which can be easily measured in modern experiments. We have shown that the anisotropy of the $2p\to 1s$ radiation and, hence, the alignment of the $2p$ state are rather sensitive to the OAM projection of the Laguerre–Gaussian beam. This sensitivity may open up a promising avenue for the diagnostics of vortex ions in accelerator and storage ring experiments. To fully explore the potential of Coulomb excitation for the ion diagnostics, however, additional analysis is needed for even wider wave packets, $\sigma_{\perp}>$ 1 $\mu$m, and to elucidate the role of their longitudinal size $\sigma_{z}$. Such an analysis, being computationally demanding and going beyond the scope of the present proof–of–principle study, will be presented in the forthcoming publications.
@@
The publication is funded by the Open Access Publishing Fund of GSI Helmholtzzentrum für Schwerionenforschung. A.S. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967.
@@
$\Psi^{LG}({\bf R},t)=\sqrt{\frac{n!}{(n+|\ell_{OAM}|)!}}\,\frac{1}{\pi^{3/4}}\left(\frac{R_{\perp}}{\sigma_{\perp}(t)}\right)^{|\ell_{OAM}|}\frac{1}{\sigma_{\perp}(t)\sqrt{\sigma_{z}(t)}}\,L_{n}^{|\ell_{OAM}|}\left(\frac{R_{\perp}^{2}}{\sigma_{\perp}^{2}(t)}\right)$
×$\times\exp\Big{\{}-\frac{i}{\hbar}t\frac{\langle p\rangle^{2}}{2m_{p}}+\frac{i}{\hbar}\langle p\rangle R_{z}+\frac{i}{\hbar}\ell_{OAM}\phi_{r}-i(2n+|\ell_{OAM}|+1)\arctan(t/t_{d,\perp})$
$-\frac{i}{2}\arctan(t/t_{d,z})-\frac{R_{\perp}^{2}}{2\sigma_{\perp}^{2}(t)}\,(1-it/t_{d,\perp})-\frac{(R_{z}-v_{p}t)^{2}}{2\sigma_{z}^{2}(t)}\,(1-it/t_{d,z})\Big{\}}\,,$ (15)
@@
The set of solutions (Appendix: wave function of the twisted proton) for $n=0,1,2...$ and $l_{OAM}=0,1,2...$ is complete and orthogonal. It describes, moreover, free rectilinear motion on average,
@@
$t_{d,{\perp}}=\frac{m_{p}\hbar}{(\sigma^{p}_{\perp})^{2}}=\frac{m_{p}\sigma^{2}_{\perp}(t_{0})}{\hbar}$ (19)
is the time after which $\sigma_{\perp}(t_{d})=\sqrt{2}\sigma_{\perp}(t_{0})$, and $\sigma^{p}_{\perp}$ is the width of the wave packet in the momentum space. Similar expressions can be obtained for the longitudinal direction:
@@
$\displaystyle\sigma_{z}(t_{0})=\frac{\hbar}{\sigma^{p}_{z}},\ t_{d,z}=\frac{m_{p}\hbar}{(\sigma^{p}_{z})^{2}}={\color[rgb]{0,0,0}\frac{m_{p}\sigma^{2}_{z}(t_{0})}{\hbar}}.$ (21)
One can note that for both cases $t_{d,{\perp}}$ and $t_{d,z}$ are inversely proportional to $\hbar/m_{p}\approx 6\times 10^{-8}$ m2/s, which implies characteristic dispersion time of about $t_{d}\sim 10^{-5}$ s for sub–$\mu$m–size proton wave packets.